OpenStudy (anonymous):

Determine the slope of the graph of x3 = ln(xy) at the point (1, e).

OpenStudy (anonymous):

I am thinking one or three.

OpenStudy (jhannybean):

have you learned implicit differentiation or anything or that sort?

OpenStudy (anonymous):

yes that is what this is

geerky42 (geerky42):

You can isolate y. \[x^3 = \ln(xy)\\~\\\ln y = x^3 - \ln x\\~\\y = e^{x^3-\ln x}\\~\\y = e^{x^3}e^{-\ln x} \\~\\y = \dfrac{e^{x^3}}{x}\] You can take derivative of y, then plug in 1 to get a answer.

OpenStudy (anonymous):

So 3 would be the answer?

geerky42 (geerky42):

\(3\)? I got \(2e\) Do you understand what I did to isolate y? What do you get after taking derivative of y?